Q1: Adrienne left Montreal biking at 30 kph. Elise left 2 hours later, traveling the same road at 40 kph. At what distance does Elise overtake Adrienne?
Answer by richwmiller(17219)

gaining speed 40 -30= 10 mph
headstart 30*2 =60 miles
time to catch up 60/10=6 hours
algebraic solution
30(t+2)=40t
30t+60=40t
60=40t-30t
60=10t
6=t
distance to meeting 6*40=240 miles
As much as I like to do these problems
Submit one at a time and no similar problems.
Those two problem are the same except for the numbers